# Solvability of N-level Systems under the Rotating Wave Approximation (RWA) and Dipole Selection Rules

### v1

@@ Abstract
We analyze the dynamics of an N-level system under the Rotating Wave Approximation and dipole selection rules. Such systems can be solved straightforwardly if the Hamiltonian can be transformed into a time-independent form. The condition under which a unitary transformation can be used to render time-dependent Hamiltonians into a time-independent form, thereby making the solution, is examined. After case-by-case analysis of different four and five-level systems, we conclude that systems having only one odd or even parity level are solvable. In contrast, the others must satisfy a condition of laser detuning for solvability.
@@ I Introduction
Laser excitations and multi-level atomic systems are simple yet fundamental tools that can be applied in various fields of quantum technology. Multi-level atomic systems are used in optical atomic clocks to achieve standard timekeeping of precision up to $10^{-19}s$ [1, 2]. Rydberg atoms are a direct application of multilevel atomic systems and are very useful in quantum gate construction and quantum information processing [3]. Successful quantum gates such as the Cirac-Zoller gate [4], and the Molmer-Sorensen gate [5] are examples of the utilization of multi-level atomic systems in quantum information science. STImulated Raman Adiabatic Passage(STIRAP) uses a 3-level lambda($\lambda$) system, and it demonstrates coherent population transfer from a nearly-degenerate ground state to the other [6, 7, 8]. Quantum computing techniques using qudits(quantum states with more than two qubits) have strong correlations with multi-level atomic systems [9]. The validity of the RWA on multi-level atomic systems when exposed to the environment is also an active research area [10]. The multi-level systems are also interesting for their applications in nonlinear optics and photonics in characterizing various non-linear processes such as four-wave mixing [11, 12, 13]. A recent work from Keck et al. demonstrated how to use a control field to make the Hamiltonian time independent under the rotating wave approximation [14]. Therefore, it is worth revisiting what we know about the solutions of general N-level systems under RWA.
This research is based on a paper from Einwohner, Wong, and Garrison that studied the solvability of N-level systems using graph theory [15]. There has also been work by Fujii et al. on the exact solutions of 3-level systems [16, 17]. We ask the same question as Einwohner, Wong, and Garrison on the solvability of various N-level systems. However, we approach it from the rotating frame and a case-by-case analysis of different systems.
@@
$\hat{H}=\sum_{n=1}^{N}\hbar\omega_{n}\ket{n}\bra{n}+\hbar(\sum_{m,n=1}^{N}\frac{\Omega_{nm}}{2}e^{-i\omega_{nm}t}\ket{n}\bra{m}+c.c.)$ (2)
$H_{0}$ represents the Hamiltonian of an isolated N-level system with no interaction between the levels. The field of interest is a multi-mode electric field with the assumption that only one field mode close to the transition frequency gets coupled [15]. Under RWA, the Hamiltonian can be represented as in equation 2 where $\Omega_{nm}=\frac{1}{\hbar}\bra{n}\hat{\vec{d}}\ket{m}\cdot\vec{E}(\vec{r})$ is the Rabi frequency between the two levels and $\omega_{nm}$ is the laser frequency of the transition between two levels.
The dipole operator($\hat{\vec{d}}$) is a classical dipole $\hat{\vec{d}}=e\hat{\vec{r}}$ which neglects the relativistic effects of the atom. Since we are interested in atomic systems concerning ultracold atoms, the approximation is valid as the cold atoms do not move at relativistic speeds. We are also ignoring the field couplings with quadrupole, octupole, etc. Dipole selection rules allow us to reduce the number of transitions. As a result, the topology of the graphs representing various systems becomes non-trivial and of interest. Therefore, we look for conditions for the solvability of different topologies governing the multi-level systems. Some applications go beyond the dipole approximation in ion-trapped quantum computing [19, 20]. Such schemes using quadrupole and octupole transitions are beyond the scope of this research.
The eigenstates (orbitals) of atoms are eigenstates of the parity operator. Since the position operator($\hat{\vec{r}}$) is odd, the matrix element $\bra{n}\hat{\vec{d}}\ket{m}$ is zero unless the two levels have opposite parities. Therefore, the Hamiltonian of the system depends on how many even-parity levels the system has. This paper analyzes every 4-level system while extending the analysis to induce the conditions for the solvability of a general N-level system.
@@ II Brief Review on the Rotating Frame Transforms
Solving the Schrodinger equation of Hamiltonians with explicit time dependence is a challenging task. There are a few approaches. The most simple approach of all is transforming the time-dependent Hamiltonian into a time-independent form using a unitary matrix and its conjugate as follows.
@@
If the transformed Hamiltonian is time-independent, it can be diagonalized for exact solutions. However, in this paper, we are only interested in the question of which systems are solvable instead of what the solutions are.
@@
$U=\sum_{n=1}^{N}e^{-i\bar{\omega}_{n}t}\ket{n}\bra{n}$ | (4)
The chosen unitary matrix has a few advantages. Since the phases cancel when multiplied with the conjugate, the second term of equation 3 has no time-dependence, and we only have to worry about $\hat{U}^{\dagger}\hat{H}\hat{U}$. However, a diagonal unitary has four degrees of freedom, and we will see that the degrees of freedom vs. the number of transitions determines the solvability of a system.
Having introduced the rotating picture and the general method of the analysis, we are ready to answer the question of which systems are solvable.
III Unconditionally Solvable Systems
@@
Since $\bar{\omega}_{n}$ are the degrees of freedom of the unitary, there exists one more $\bar{\omega}_{n}$ than the transition frequencies($\omega_{mn}$). Following the same analysis on M, Y, and W systems, one sees that there also exists one more degree of freedom than the transition frequency in each 3+1 system. With the freedom to choose an extra constraint for the unitary, we conclude that every 3+1 system is unconditionally solvable.
@@
The following is the set of equations to be solved for the 4+1 $\lambda$ system.
@@
It is seen that the first three equations are identical to the equations of the 3+1 $\lambda$ system. It can be shown that all the other $4+1$ systems are merely a one-level extension to the respective $3+1$ systems. This claim can be extended to N-level systems as follows: (N-1)+1 level systems are merely (N-4) level extensions to 3+1 systems. In other words, there exist (N-4) more different levels for (N-1)+1 systems, but in such systems, the degrees of freedom of the unitary are always one above the number of transitions. Therefore, we conclude that all (N-1)+1 systems are unconditionally solvable.
IV Conditionally Solvable Systems
We have seen in the previous section that the condition for solvability depends on the degrees of freedom of the unitary matrix and the number of transitions of a system. Mathematically, for an N-level system with n number of even levels, the diagonal unitary possesses N degrees of freedom, and the number of transitions is the product of the number of even and odd levels. One would think that the condition for solvability would be $N\leq N(N-n)$, but we will see that the system is not unconditionally solvable when the degrees of freedom are the same as the number of transitions.
@@ IV Conditionally Solvable Systems
There are four degrees of freedom and four transition frequencies, so the system has the potential to be solvable. When attempted, we encounter a condition of $\omega_{13}+\omega_{34}=\omega_{12}+\omega_{23}$. The system is solvable if the condition is met, unsolvable otherwise. We define the laser detuning of the diamond system as $\Delta_{D}=\omega_{13}+\omega_{34}-\omega_{12}-\omega_{23}$, and the Hamiltonian is time-independent if the lasers are tuned in a way that the system detuning is zero.
The hourglass and the trapezium system can also be solvable if a detuning condition is met. The detuning depends on the topology of the system and different systems have different detunings. This paper will not go into detail about the detuning of each system. Considering the simplicity of 4-level matrix mechanics, we have found the detunings of the hourglass and trapezium systems as $\Delta_{T}=\omega_{14}-(\omega_{12}+\omega_{23}+\omega_{34})$ and $\Delta_{h}=\omega_{13}+\omega_{24}-(\omega_{23}+\omega_{14})$ respectively.
One final remark about the detuning is that Einwohner, Wong, and Garrison approached the same problem with graph theory and they claimed that a system is solvable if the way to go around the graph is unchanged [15]. For the diamond system, the detuning that we define is merely the difference between the way up from the left and the way up from the right. This pattern reoccurs for other 2+2 systems as well and therefore, our results agree with previous work.
@@ IV.2 Higher Level Systems
We have seen that (N-1)+1 level systems are solvable since they meet the condition of solvability: the degrees of freedom must be less than the number of transitions. For an N-level system with an even level, the degree of freedom is N, and the number of transitions is n(N-n). The only possible way of $N<n(N-n)$ is if $n=1$. Otherwise, we will have more equations than the degree of freedom of the unitary, and the system will not be unconditionally solvable.
It is also evident that the higher the number of transitions is, the more equations are left unsolved, and the transformed Hamiltonian will have leftover detuning-dependent phases. In other words, higher-level systems will have more detuning conditions upon the first rotation. We can use the same form of unitary (equation 4) to further rotate $\hat{H}^{\prime}$, and we see that $\hat{H}^{\prime}$ can be reduced into only one detuning-dependent term eventually. In other words, the Hamiltonian of every system except the unconditionally solvable ones, can be reduced upon multiple rotations into the Hamiltonian with only one detuning-dependent phase. It is impossible to further reduce the time dependence of the Hamiltonian from one detuning-dependent time-dependent phase. The intuitive justification is that when one of the entries of the Hamiltonian and its complex conjugate are the only time-dependent terms, there exists no other frame that can remove the time dependence.
@@ V Discussion and Conclusion
In this paper, we have revisited the solvability of N-level systems under the RWA. Three-level $\lambda$ systems are used in STImulated Raman Adiabatic Passage(STIRAP) for information transfer between levels [6]. Moreover, STIRAP can transfer the population from the ground state (|$\ket{g_{1}}$ to a nearly-degenerate ground state with higher energy (|$\ket{g_{2}}$) without coherence. A universal X-gate can be constructed if the reverse operation that transfers the population from |$\ket{g_{2}}$ to |$\ket{g_{1}}$ is implemented. We are interested in whether we can utilize higher-level systems using adiabatic transitions to construct a universal quantum logic gate. Since $(N-1)+1$ systems are unconditionally solvable under RWA, we plan to analyze such systems for coherence dynamics.
Moreover, the detuning conditions that will make a system solvable are interesting, and they should be explored as well. We have done case by case analysis up to five-level systems and it appears that all unsolvable systems can be reduced into Hamiltonians containing one detuning-dependent time-evolving phase. However, we lack concrete mathematical proof to support this claim and the graph theory seems to be the approach.
Nevertheless, we have seen the unconditionally solvable systems under RWA, and the solutions should be explored as it may be useful in constructing a quantum universal logic gate. The solvability conditions for the other systems are worth exploring as we claim that every N-level system under the RWA is solvable if the lasers are detuned according to the system’s topology. Solving the multi-level atomic systems without the RWA is another interesting research question. This research only focused on the scenarios where the RWA is valid. To approach the Hamiltonians with the counter-rotating term, one can attempt using the affective Hamiltonian theory, potentially to calculate the Bloch-Siegert shift of different systems [21].
In this research, we have reviewed N-level systems under the RWA by a case-by-case analysis in the rotating frame. Although this approach is not the only way to solve a time-dependent Hamiltonian, it is the simplest and analytic approach to see which systems can be solved analytically.

### v2

# Conditions for Time-Independence of N-level Systems under the Rotating Wave Approximation (RWA) and Dipole Selection Rules

@@ Abstract
We analyze the time-dependence of N-level systems under the Rotating Wave Approximation and dipole selection rules. Such systems can be solved straightforwardly if the Hamiltonian can be transformed into a time-independent form. The conditions under which a unitary transformation can be used to render time-dependent Hamiltonians into a time-independent form, thereby making the solution, are examined. After case-by-case analysis of different four and five-level systems, we conclude that systems having only one odd or even parity level achieve time-independence. In contrast, the others must satisfy a condition of laser detuning to achieve time-independence.
Rotating Wave Approximation, N-level systems, Dipole selection rules, Hamiltonian time-independence, Rotating frame transform, System detuning conditions
@@
This is the author-accepted version of the manuscript that has been accepted for publication in Journal of Modern Optics (Taylor & Francis, 2025).
@@ I Introduction
Laser excitations and multi-level atomic systems are simple yet fundamental tools that can be applied in various fields of quantum technology. Multi-level atomic systems are used in optical atomic clocks to achieve standard timekeeping of precision up to $10^{-19}s$ [1, 2]. Rydberg atoms are a direct application of multilevel atomic systems and are very useful in quantum gate construction and quantum information processing [3]. Successful quantum gates such as the Cirac-Zoller gate [4], and the Molmer-Sorensen gate [5] are examples of the utilization of multi-level atomic systems in quantum information science. STImulated Raman Adiabatic Passage(STIRAP) uses a 3-level lambda($\lambda$) system, and it demonstrates coherent population transfer from a nearly-degenerate ground state to the other [6, 7, 8]. Quantum computing techniques using qudits(quantum states with more than two qubits) strongly correlate with multi-level atomic systems [9]. The validity of the RWA on multi-level atomic systems when exposed to the environment is also an active research area [10]. The multi-level systems are also interesting for their applications in nonlinear optics and photonics in characterizing various nonlinear processes such as four-wave mixing [11, 12, 13]. A recent work from Keck et al. demonstrated how to use a control field to make the Hamiltonian time-independent under the rotating wave approximation [14]. Beyond the RWA, counter-rotating terms can introduce significant modifications to level dynamics, such as Bloch–Siegert shifts and altered transition pathways [15]. Recent work by Bugarth et al. and Paing explores such effects in two-level contexts, using the Jaynes-Cummings model while highlighting issues with the RWA [16, 17]. It can be seen that although the RWA has theoretical inconsistencies and limitations, analytical calculations beyond the RWA is challenging even for a simple two-level model of an atom. It is beyond the scope of the research to investigate non RWA contributions on higher-level systems. Therefore, it is worth revisiting what we know about the solutions of general N-level systems under RWA.
This research is based on a paper from Einwohner, Wong, and Garrison that studied the solvability of N-level systems using graph theory [18]. There has also been work by Fujii et al. on the exact solutions of 3-level systems [19, 20]. We ask a similar question as Einwohner, Wong, and Garrison on the condition for time-independence of various N-level systems. However, we approach it from the rotating frame and a case-by-case analysis of different systems.
@@
[…]
The dipole operator($\hat{\vec{d}}$) is a classical dipole $\hat{\vec{d}}=e\hat{\vec{r}}$ which neglects the relativistic effects of the atom. Since we are interested in atomic systems concerning ultracold atoms, the approximation is valid as the cold atoms do not move at relativistic speeds. We are also ignoring the field couplings with quadrupole, octupole, etc. Dipole selection rules allow us to reduce the number of transitions. As a result, the topology of the graphs representing various systems becomes non-trivial and of interest. Therefore, we look for conditions for complete time-independence of different topologies governing the multi-level systems. Some applications go beyond the dipole approximation in ion-trapped quantum computing [22, 23]. Such schemes using quadrupole and octupole transitions are beyond the scope of this research.
[…]
@@ II Brief Review on the Rotating Frame Transforms
Solving the Schrodinger equation of Hamiltonians with explicit time dependence is a challenging task. There are a few approaches. The simplest approach of all is transforming the time-dependent Hamiltonian into a time-independent form using a unitary matrix and its conjugate as follows.
@@
If the transformed Hamiltonian is time-independent, it can be diagonalized for exact solutions. However, in this paper, we are only interested in the question of which systems can be rendered into a time-independent form instead of what the solutions are.
@@
$|U=\sum_{n=1}^{N}e^{-i\bar{\omega}_{n}t}\ket{n}\bra{n}$ (4)
The chosen unitary matrix has a few advantages. Since the phases cancel when multiplied with the conjugate, the second term of equation 3 has no time-dependence, and we only have to worry about $\hat{U}^{\dagger}\hat{H}\hat{U}$. However, a diagonal unitary has four degrees of freedom, and we will see that the degrees of freedom vs. the number of transitions determines whether a system can be transformed into a time-independent form or not.
Having introduced the rotating picture and the general method of the analysis, we are ready to answer the question of which systems are time-independent.
III Unconditionally Time-Independent Systems
@@
Since $\bar{\omega}_{n}$ are the degrees of freedom of the unitary, there exists one more $\bar{\omega}_{n}$ than the transition frequencies($\omega_{mn}$). Following the same analysis on M, Y, and W systems, one sees that there also exists one more degree of freedom than the transition frequency in each 3+1 system. With the freedom to choose an extra constraint for the unitary, we conclude that every 3+1 system can be transformed into a time-independent Hamiltonian without the need for any detuning conditions.
@@
The following is the set of equations to be solved for the time-independence of 4+1 $\lambda$ system.
@@
It is seen that the first three equations are identical to the equations of the 3+1 $\lambda$ system. It can be shown that all the other $4+1$ systems are merely a one-level extension to the respective $3+1$ systems. This claim can be extended to N-level systems as follows: (N-1)+1 level systems are merely (N-4) level extensions to 3+1 systems. In other words, there exist (N-4) more different levels for (N-1)+1 systems, but in such systems, the degrees of freedom of the unitary are always one above the number of transitions. Therefore, we conclude that all (N-1)+1 systems can unconditionally be transformed into a time-independent Hamiltonian.
IV Conditionally Time-Independent Systems
We have seen in the previous section that the condition for time-independence depends on the degrees of freedom of the unitary matrix and the number of transitions of a system. For an N-level system with n number of even levels, the diagonal unitary possesses N degrees of freedom, and the number of transitions is the product of the number of even and odd levels. One would think that the condition for time-independence would be $N\leq N(N-n)$, but we will see that the system requires a detuning condition for complete time-independence when the degrees of freedom are the same as the number of transitions.
@@ IV Conditionally Time-Independent Systems
There are four degrees of freedom and four transition frequencies, so the system has the potential to be time-independent. When attempted, we encounter a condition of $\omega_{13}+\omega_{34}=\omega_{12}+\omega_{23}$. The system can be transformed into a time-independent form if the condition is met. We define the system detuning of the diamond system as $\Delta_{D}=\omega_{13}+\omega_{34}-\omega_{12}-\omega_{23}$, and the Hamiltonian is time-independent if the lasers are tuned in a way that the system detuning is zero.
The hourglass and the trapezium system can also be transformed into complete time-independence if a detuning condition is met. The detuning depends on the topology of the system, and different systems have different system detuning conditions. This paper will not go into detail about the detuning of each system. Considering the simplicity of 4-level matrix mechanics, we have found the detunings of the hourglass and trapezium systems as $\Delta_{T}=\omega_{14}-(\omega_{12}+\omega_{23}+\omega_{34})$ and $\Delta_{H}=\omega_{13}+\omega_{24}-(\omega_{23}+\omega_{14})$ respectively.
One final remark about the detuning is that Einwohner, Wong, and Garrison approached the same problem with graph theory, and they claimed that a system is solvable if the way to go around the graph is unchanged [18]. For the diamond system, the detuning that we define is merely the difference between the way up from the left and the way up from the right. This pattern reoccurs for other 2+2 systems as well and therefore, our results agree with previous work.
@@ IV.2 Higher Level Systems
We have seen that (N-1)+1 level systems are time-independent since they meet the condition of time-independence: the degrees of freedom must be less than the number of transitions. For an N-level system with one even level, the degree of freedom is N, and the number of transitions is n(N-n). The only possible way of N<n(Nn)N<n(N-n) is if $n=1$. Otherwise, we will have more equations than the degree of freedom of the unitary, and the system will remain time-dependent.
It is also evident that the higher the number of transitions is, the more equations are left unsolved, and the transformed Hamiltonian will have leftover detuning-dependent phases. Higher-level systems will have more detuning conditions upon the first rotation. We can use the same form of unitary (equation 4) to further rotate $\hat{H}^{\prime}$, and we see that $\hat{H}^{\prime}$ can be reduced into only one detuning-dependent term eventually. In other words, the Hamiltonian of every system except the unconditionally time-independent ones, can be reduced upon multiple rotations into the Hamiltonian with only one detuning-dependent phase. It is impossible to further reduce the time dependence of the Hamiltonian from one detuning-dependent phase. The intuitive justification is that when one of the entries of the Hamiltonian and its complex conjugate are the only time-dependent terms, there exists no other frame that can remove the time-dependence.
@@ V Discussion and Conclusion
[…]
Moreover, the search for detuning conditions that will make a system completely time-independent remains an open area of research. We have done case by case analysis up to five-level systems, and it appears that all systems with detuning conditions can be reduced into Hamiltonians containing one detuning-dependent time-evolving phase. However, we lack concrete mathematical proof to support this claim and the graph theory seems to be the approach.
Nevertheless, we have seen the unconditionally time-independent systems under RWA, and the solutions should be explored, as it may be useful in constructing a quantum universal logic gate. The detuning conditions for the other systems are worth exploring, as we claim that every N-level system under the RWA is time-independent if the lasers are detuned according to the detuning conditions that depend on the topology of the system.
The present analysis focuses on conditions in which a multilevel atomic system can be transformed into a time-independent form under the RWA. Going beyond RWA produces additional frequency shifts, such as the Bloch-Siegert shift, which can alter the detuning conditions identified here. In some topologies, the introduction of such counter-rotating terms may prevent the complete removal of time-dependence.
Extending the present classification to include such effects lies beyond the scope of this article because the calculations become significantly more involved even for a simple two-level system. In our recent work, we applied the Magnus expansion to the Jaynes–Cummings model without the RWA and found that the second-order terms already yield beyond the RWA phenomena, predicting not only the Bloch–Siegert shift but also an atom-induced field-squeezing interaction [17]. The fact that such a nontrivial structure emerges in the minimal two-level case shows how challenging a full general N-level treatment would be.
Nevertheless, recent studies such as Bugarth et al. [16] and our Magnus-based analysis suggest that effective Hamiltonian and perturbative approaches could be adapted to general N-level systems to quantify beyond-RWA corrections to the time-independence criteria established here. Pursuing this extension would connect the present classification directly to experimentally observable beyond-RWA phenomena and remains an interesting challenge for future work.
In this research, we have reviewed N-level systems under the RWA by a case-by-case analysis in the rotating frame. Although this approach is not the only way to solve a time-dependent Hamiltonian, it is the simplest and analytic approach to see which systems have a Hamiltonian that can be exponentiated for time-evolution.